# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v3

@@ 4.2 Study area
After identifying and compiling the BM near-miss events for each observation unit (segments and intersections), the frequencies of extreme V–V and V–I events were systematically aggregated and visualized in Fig. 4.2. This plot illustrates the spatial distribution of extreme near-miss frequencies across 28 locations, comprising 10 intersections and 18 segments, organized into intersection zones, primary-direction segments, and opposite-direction segments. This spatial aggregation confirms the integrity of the BM dataset and provides the foundation for modeling. The study extends prior approaches by incorporating both V–V and V–I near misses into the HBSGRP–UGEV framework, explicitly integrating roadway geometry and infrastructure features alongside vehicle dynamics. Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced. For example, high steering angles, often indicating abrupt lateral maneuvers, were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021). Fig. 4.2 highlights several notable spatial patterns. Opposite-direction segments (S82, S72, S62) exhibited the highest V–I frequencies, suggesting greater lateral variability and exposure to curb or median near-misses. Primary-direction segments (S11, S21, S81) showed elevated V–V near-miss frequencies, consistent with midblock environments dominated by longitudinal interactions such as close-following, lane changes, or overtaking. Among intersections, I7, I8, and I9 emerged as critical hotspots, with elevated frequencies of both near-miss types. This dual concentration reflects the operational complexity of intersections, where near-missing maneuvers, pedestrian flows, and signal control interact to amplify crash potential.
The distribution and magnitude of near-miss counts across all locations validate the effectiveness of the BM approach in capturing rare but safety-critical events. Specifically, nearly all study sites meet or exceed the recommended minimum threshold of 30 BM events for reliable GEV-based inference (Zheng et al., 2014), with the exception of intersection I2. This ensures that the dataset is statistically robust for subsequent HBSGRP–UGEV analysis. The spatial heterogeneity revealed in Fig. 4.2 further underscores the need to incorporate direction-specific geometric and behavioral covariates within the modeling framework. These results provide compelling justification for developing localized, high-resolution safety models that integrate both roadway configuration and observed vehicle dynamics, thereby enabling precise and context-sensitive real-time risk assessments in complex urban environments.
@@ 5 Modeling estimation result
This study develops a series of Bayesian hierarchical UGEV models to estimate corridor-wide COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami. Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction. The theoretical basis and implementation of this sampling methodology are detailed in prior work (2025b). Each V–V and V–I interaction is treated as an independent block. Since extreme events are relatively sparse at individual sites, direct location-specific estimation often yields unstable posteriors with wide uncertainty intervals. To address this issue, a corridor-wide grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while preserving local heterogeneity.
@@
The HBSGRP–UGEV model is the primary analytical framework, with a simpler fixed-effect hierarchical model (HBSFP) also estimated for comparison. For each near-miss type (V–V and V–I), two models were therefore developed: (1) HBSFP and (2) HBSGRP. The grouped random parameter formulation in HBSGRP allows selected coefficients to vary across spatial clusters, thereby capturing latent heterogeneity associated with interaction dynamics and roadway characteristics. Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), with lower values indicating superior predictive fit (Zheng et al., 2019).
@@
Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference. Its support for block updating and parallel execution makes it well-suited to modeling spatial dependence, random effects, and extreme value behavior within complex data structures. Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation; two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference. Convergence was assessed both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, confirming reliable convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009). This procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
@@
The modeling framework incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the nonstationary UGEV distribution to capture contextual and behavioral variability in COR. For both V–V and V–I interactions, dynamic covariates included relative (speed, distance, acceleration, deceleration, headway, steering angle), jerk, and traffic volume, thereby capturing multiple dimensions of vehicle dynamics. Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position of vehicle.
@@ 5.1 Posterior results and model performance
Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments. Across all cases, the HBSGRP structure consistently outperformed the HBSFP baseline model, confirming its capacity to capture unobserved heterogeneity and location-level variability. The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments. These correspond to relative reductions of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts. Such improvements exceed the widely accepted threshold of five DIC units, indicating substantively better model fit (El-Basyouny and Sayed, 2012b, a). To further evaluate robustness, WAIC and LOOIC were also computed, both of which produced rankings consistent with DIC, thereby reinforcing the superior performance of the HBSGRP. By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion, a capability absent in fixed-effect structures. Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
@@ 5.1.1 Intersection-level
The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions from V-V interactions at signalized intersections. In the modeling stage, all covariates were considered as random effects; however, roadway characteristics and few vehicle dynamics were exhibited negligible variation across sites and did not improve posterior fit. To maintain parsimony and interpretability, random effects were restricted to the intercept and two dynamic covariates (relative speed and distance), while other covariates were retained as fixed effects, following recommendations from (Anis et al., 2025a). This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence. The resulting specification captures both site-specific dynamics and broader geometric constraints, group-specific random effects illustrated in Fig. 5.1.1 with posterior summaries reported in Table 5.
@@
    Hierarchical Bayesian Spatial Fixed Parameter
@@
    Hierarchical Bayesian Spatial Grouped Random Parameter
@@
Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, highlighting the added value of allowing speed and spacing effects to vary across intersections. The direction and magnitude of these coefficients provide insights into the mechanisms underlying COR. Relative speed and distance emerge as the dominant dynamics, influencing both the location ($\mu$) and scale ($\sigma$) parameters. For $\mu$, relative speed has a positive effect, indicating that larger speed differentials between vehicles sharply reduce 2D-TTC and intensify near-miss severity. In contrast, relative distance shows a negative effect, meaning shorter headways systematically elevate severity by limiting available buffer space for evasive action. For $\sigma$, both covariates carry negative signs, implying that high speeds and short distances compress variability, yielding consistently critical outcomes with little room for driver recovery through braking or swerving. Taken together, these results confirm that speed amplifies severity while short headways exacerbate risk and reduce variability, producing deterministic near-miss outcomes under constrained conditions. Fig. 5.1.1 illustrates the spatial heterogeneity across intersections. While most locations show consistent trends, the I2 site exhibits unusually wide credible intervals for the $\mu$ effects and intervals exceeding zero for the $\sigma$ across the intercept, relative speed, and relative distance. This instability is attributable to I2’s small BM sample size, which weakens posterior precision and increases uncertainty in group-level effects. Despite localized uncertainty, the dominant intersection-level trend persists, with relative speed and distance exerting stable and significant effects on near-miss outcomes.
Beyond random effects, several covariates were modeled as fixed parameters due to their stable contributions across intersections. These effects can be interpreted in terms of driver behavior. These factors highlight the roles of longitudinal dynamics, lateral control, maneuver type, and exposure in shaping near-miss outcomes. Table 5 shows that relative acceleration has a significant negative effect on both the location and scale parameters, suggesting smoother and less variable near misses when vehicles accelerate through green phases or clear intersections without abrupt interaction. By contrast, relative deceleration increases severity ($\mu$) but reduces variability ($\sigma$), capturing emergency braking events. These situations sharply intensify near misses but also produce uniform outcomes, since drivers in crisis have limited response options. Jerk, representing the rate of change in acceleration, shows the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$. This suggests that rapid throttle–brake oscillations, often signs of hesitation or overcorrection, slightly lower average severity but introduce greater unpredictability, underscoring the instability of abrupt corrective maneuvers.
Lateral vehicle dynamics also play an important role in near-miss risk estimation. Steering difference is a significant factor in the HBSGRP model, showing a positive association with $\mu$ and a negative association with $\sigma$. This indicates that abrupt steering maneuvers intensify near-miss severity but produce more consistent outcomes within constrained intersection geometries. Heading difference shows a similar pattern, with a positive effect on $\mu$ and a negative effect on $\sigma$. This suggests that even small misalignments in vehicle trajectory elevate severity while reducing variability, reflecting the rigid geometric constraints of signalized approaches.
The categorical indicator for left-turn movements is significant, with positive effects on both $\mu$ and $\sigma$. This finding is consistent with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies. Left turns therefore not only elevate near-miss severity but also increase variability, reflecting differences in intersection design and driver behavior under turning conditions. By contrast, traffic volume is statistically insignificant for both parameters but was retained in the model to represent exposure. Its lack of effect suggests that short-duration extreme 2D-TTC events are driven more by instantaneous vehicle interactions than by aggregate demand.
@@
The HBSGRP model for V–I near misses specified the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in severity and variability across intersections. Relative speed (in $\mu$) and relative distance (in $\sigma$) were also modeled as random effects (Fig. 5.1.1), although their 95% credible intervals overlapped zero. They were nevertheless retained on theoretical grounds, as their inclusion improved overall model fit. This specification reflects behavioral intuition: heterogeneity in speed effects is more pronounced in severity (location), whereas heterogeneity in clearance effects is more evident in variability (scale). All other covariates were treated as fixed effects, reflecting their comparatively stable behavioral and geometric influences. Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC relative to the HBSFP specification, suggesting that the grouped random parameter structure adds explanatory nuance but produces smaller performance gains than in the V–V case.
@@
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.147 0.056 $0.285,-0.027]^{\dagger}$g -0.177 0.056 $0.289,-0.047]^{\dagger}$
@@
    Hierarchical Bayesian Spatial Fixed Parameter
@@
    Hierarchical Bayesian Spatial Grouped Random Parameter
@@
In the location parameter, several interpretable effects emerge. Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC and elevates severity. Relative deceleration also enters with a significant negative sign, contrasting with its positive effect in V–V near misses. This suggests that braking near infrastructure often reflects proactive yielding or evasive maneuvers, which mitigate severity rather than intensify it. Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity. Left-turn movement, by contrast, shows a significant negative association, suggesting that left-turn trajectories provide greater clearance to boundaries and reduce severity compared with right turns. This pattern may reflect geometric characteristics such as wider turning radii or reduced overlap with curbs and crosswalks during left-turn phasing. Finally, relative speed—although modeled as a random effect—does not reach significance in the pooled analysis, reinforcing that V–I severity is less sensitive to speed differentials than V–V encounters.
The scale parameter results provide additional insight into variability. The negative random intercept indicates baseline differences across intersections, with some sites producing more predictable outcomes than others. Relative speed shows a positive and significant association with $\sigma$, implying that faster approaches to infrastructure increase variability in severity—in contrast to the V–V case, where speed compressed dispersion. This divergence reflects the diverse ways drivers respond to boundary proximity: some brake aggressively, while others maintain speed or adjust trajectory. Relative acceleration is also positive and significant, reinforcing that heterogeneous acceleration behavior near boundaries broadens outcome variability. By contrast, relative deceleration has a significant negative effect, indicating that decisive braking compresses the distribution into consistently critical outcomes. Jerk is similarly positive, confirming that abrupt throttle–brake oscillations amplify variability. Lateral dynamics play the opposite role: heading and steering differences both reduce variability, suggesting that once a lateral maneuver is committed, outcomes become more predictable within intersection geometry. Finally, left-turn movement increases variability despite lowering mean severity, consistent with heterogeneous layouts and driver strategies in left-turn contexts.
@@
The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail. This indicates a physical and behavioral ceiling on severity, constrained by roadway geometry and driver response capacity. Spatial variation in $\xi$ further underscores the influence of intersection design and infrastructure placement on the extremal properties of near-miss distributions. Taken together, the intersection-level results suggest that V–I near misses are shaped more by clearance, braking, and lateral maneuvers than by speed differentials. In contrast, V–V near misses are dominated by high speeds, which produce deterministic severity, whereas V–I outcomes exhibit greater variability driven by the interplay of infrastructure geometry, driver avoidance strategies, and intersection control.

### v4

@@ 4.4 Study area
After extracting block extremes per valid interaction window, the BM extremes are aggregated by location and summarized in Fig. 4.4. The figure reports extreme near-miss frequencies across 28 locations. To contextualize exposure, Tables 34 also report frame-level volumes, which varied across scenarios. Because the ego-centric LiDAR has a limited sensing range, the observable traffic volume varies across frames and interaction windows and is therefore used as a local exposure measure to contextualize 2DTTC-based extremes. The study extends prior EVT-based near-miss work by jointly modeling V–V and V–I extremes under the HBSGRP–UGEV framework and incorporating roadway geometry and infrastructure context alongside vehicle dynamics. In addition to kinematic predictors emphasized in earlier studies (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015), we include geometric covariates, such as lane configuration and boundary context; for example, steering-related measures capture abrupt lateral maneuvers associated with boundary incursions (Gilbert et al., 2021). Fig. 4.4 also reveals clear spatial patterns: opposite-direction segments (S82, S72, S62) show the highest V–I extreme frequencies, while primary-direction segments (S11, S21, S81) exhibit elevated V–V frequencies consistent with midblock longitudinal interactions. Among intersections, I7–I9 emerge as hotspots for both near-miss types, reflecting the compounded operational complexity of signal control, turning movements, and interaction density.
Overall, the distribution of block extremes supports the BM approach in capturing rare but safety-critical events. Nearly all sites meet or exceed the commonly cited minimum of 30 observations for GEV-based inference (Zheng et al., 2014), except for I2, which provides a statistically adequate basis for HBSGRP–UGEV estimation. The observed heterogeneity across directions and facility types further motivates the inclusion of direction and location-specific effects to enable context-sensitive, corridor-scale crash-risk estimation.
@@ 5 Modeling estimation result
This study develops a set of Bayesian hierarchical UGEV models to estimate corridor-wide crash occurrence risk (COR) from extreme 2D-TTC near-miss events along Biscayne Boulevard in Miami. Extreme near-misses are extracted using a BM strategy; each V–V or V–I defines a block extreme used for EVT inference. Because extreme events are sparse at individual sites, direct site-by-site estimation can yield unstable posteriors and wide credible intervals. To improve statistical stability while preserving spatial structure, we adopt a corridor-wide grouped formulation. Each site defines one group $k$, and all block extremes observed at that site are assigned to the corresponding group. This enables partial pooling through corridor-level hyperparameters while allowing selected effects to vary across the 28 groups.
@@
The HBSGRP–UGEV model is the primary analytical framework, with a simpler fixed-effect hierarchical model (HBSFP) also estimated for comparison. For each near-miss type (V–V and V–I), two models were therefore developed: (1) HBSFP and (2) HBSGRP. The grouped random-parameters formulation in HBSGRP allows selected coefficients to vary across spatial clusters, thereby capturing latent heterogeneity in interaction dynamics and roadway characteristics. To compare alternative model specifications, we relied on several Bayesian information criteria commonly used in roadway safety research. The Deviance Information Criterion (DIC) (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a) is frequently reported as a summary measure of model fit, with lower values indicating superior predictive fit (Zheng et al., 2019). However, prior studies have demonstrated that its numerical value can be affected by the model’s parameterization (e.g., Geedipally et al., 2012, 2014). Align with earlier safety studies, a difference greater than five DIC units was taken to indicate a meaningfully better fit between competing models (El-Basyouny and Sayed, 2012b, a). To strengthen the assessment of predictive performance and reduce sensitivity to parameterization, additionally report the Widely Applicable Information Criterion (WAIC) and the Leave-One-Out Cross-validation Information Criterion (LOOIC), which have been recommended as more robust alternatives to DIC for Bayesian model comparison (Vehtari et al., 2017) and have seen growing use in recent roadway safety applications (Dzinyela et al., 2024; Islam et al., 2023; Khodadadi et al., 2023).
@@
Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference. Its support for block updates and parallel execution makes it well-suited for fitting hierarchical models with grouped random effects and extreme-value behavior within complex data structures. Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation; two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference. Convergence was assessed both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, confirming reliable convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009). This procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
@@
The modeling framework incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the nonstationary UGEV distribution to capture contextual and behavioral variability in COR. For both V–V and V–I interactions, dynamic covariates included relative (speed, distance, acceleration, deceleration, headway, steering angle), jerk, and traffic volume, thereby capturing multiple dimensions of vehicle dynamics. Geometric and infrastructure-related covariates were also included: lane number, lane width, driveway intensity, median type, and vehicle lane position.
@@ 5.1 Posterior results and model performance
Tables 58 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-miss events at intersections and segments. Across all cases, the HBSGRP structure consistently outperforms the HBSFP baseline, indicating a better ability to capture unobserved heterogeneity and location-level variability. For the HBSGRP models, the DIC values are 508.9 (V–V) and 3179.6 (V–I) at intersections and 4189.0 (V–V) and 6550.0 (V–I) at segments. These correspond to relative reductions of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts. Such improvements exceed the widely accepted threshold of five DIC units, indicating substantively better model fit (El-Basyouny and Sayed, 2012b, a). To further evaluate robustness, WAIC and LOOIC were also computed, both of which produced rankings consistent with DIC, thereby reinforcing the HBSGRP’s superior performance. By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion, a capability absent in fixed-effect structures. Collectively, these results demonstrate that grouped random parameters models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
@@ 5.1.1 Intersection-level
The intersection-level HBSGRP–UGEV model was designed to capture heterogeneity in extreme 2D-TTC from V–V interactions at signalized intersections. In an initial specification, all covariates were allowed to vary across site groups; however, the final random-parameter set was determined using a theory-first rationale and then empirically confirmed and parsimonious. Specifically, random effects were retained for the intercept and two varying-dynamics covariates, relative (speed and distance). These quantities directly govern the short-horizon evolution of 2D-TTC and are expected to vary across intersections with signal phasing, approach geometry, queue discharge conditions, and driver compliance. In contrast, other vehicle dynamics (e.g., relative acceleration/deceleration, jerk, heading/steering differences) are typically more episodic and noise-sensitive at high frequency; once closing dynamics are accounted for, their between-site variation was not consistently identifiable and contributed limited additional explanatory power.
To avoid overfitting and preserve interpretability, restricted random effects to this minimal, theoretically motivated subset and relied on hierarchical partial pooling to shrink site-specific coefficients toward corridor-level means when information was limited. Candidate random-effects structures were compared using posterior stability (non-degenerate random-effect variances), predictive information criteria, and posterior predictive checks; covariates with random-effect variances concentrated near zero or unstable posteriors were retained as fixed effects (Anis et al., 2025a). The resulting specification captures intersection-to-intersection variability in closing behavior while maintaining a parsimonious, generalizable structure; group-specific random effects are illustrated in Fig. 5.1.1, with posterior summaries reported in Table 5.
@@
    Hierarchical Bayesian Structure Fixed Parameter
@@
    Hierarchical Bayesian Structure Grouped Random Parameters
@@
Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, highlighting the added value of allowing speed and spacing effects to vary across intersections. The direction and magnitude of these coefficients provide insights into the mechanisms underlying COR. Relative speed and distance emerge as the dominant dynamics, influencing both $\mu$ and $\sigma$ parameters. Relative speed exhibits a positive association with the location component and a negative association with the scale component (Fig. 8(b), Fig. 8(e)), whereas relative distance exhibits a negative association with the location component and a negative association with the scale component (Fig. 8(c), Fig. 8(f)). Taken together, these changes imply that higher closing-rate conditions and tighter spacing shift the fitted extreme distribution toward more critical tail behavior while simultaneously concentrating the distribution. Fig. 5.1.1 illustrates the spatial heterogeneity across intersections, and the corresponding group-level posterior intervals are provided. While most locations show consistent trends, site I2 exhibits wider credible intervals for the location effects and less stable scale-related posteriors, attributable to its smaller BM sample size and the resulting weaker posterior precision.
Beyond random effects, several covariates were modeled as fixed effects because of their stable contributions across intersections. These effects can be interpreted in terms of driver behavior. These factors highlight the roles of longitudinal dynamics, lateral control, maneuver type, and exposure in shaping near-miss outcomes. Table 5 shows that relative acceleration is negatively associated with both the location and scale components, implying lower predicted exceedance probability and a more concentrated extreme distribution under higher-acceleration conditions. By contrast, relative deceleration is positively associated with the location component and negatively associated with the scale component, implying higher predicted exceedance probability with a more concentrated tail under braking-dominant interactions. Jerk is negatively associated with the location component but positively associated with the scale component, implying lower central extreme levels but greater dispersion in tail behavior, i.e., more variability in predicted exceedance probability across comparable contexts.
Lateral control variables also contribute meaningfully at intersections. Steering difference shows a positive association with the location component and a negative association with the scale component, while heading difference exhibits the same sign pattern. Taken together, these coupled changes imply that larger lateral corrections and trajectory misalignment shift the fitted extreme distribution toward more crash-prone tail behavior while concentrating it, resulting in higher exceedance probability under unstable lateral maneuvers in constrained intersection geometry. The left-turn indicator is significant, with positive associations in both the location and scale components, indicating that left-turn maneuvers shift the extreme 2D-TTC distribution toward more crash-prone behavior while increasing dispersion, consistent with the added conflict complexity of opposing flows and heterogeneous gap-acceptance. Traffic volume is not statistically significant in either component; it is retained as an exposure control, and its weak effect suggests that block-level extremes are dominated by instantaneous interaction states rather than by aggregate demand within the short scenario windows.
@@
The intersection-level HBSGRP model for V–I near-misses specified the intercept as a grouped random parameters in both the location and scale components (Fig. 5.1.1) to capture baseline heterogeneity across intersections. Relative speed was modeled as a grouped random effect in the location component, and relative distance was modeled as a grouped random effect in the scale component. Although these effects are weakly identified at some sites (credible intervals include zero), they were retained because they are mechanistically relevant to boundary-encroachment dynamics and because they improve predictive fit relative to the fixed-parameter specification. All remaining covariates were treated as fixed effects. Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC versus HBSFP, indicating smaller gains from grouped random parameters for V–I than for V–V at intersections.
@@
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.147 0.056 $0.285,-0.027]^{\dagger} -$0.177 0.056 $0.289,-0.047]^{\dagger}$
@@
    Hierarchical Bayesian Structure Fixed Parameter
@@
    Hierarchical Bayesian Structure Grouped Random Parameters
@@
Relative deceleration shows a negative association with the location parameter and a negative association with the scale parameter, indicating that stronger braking near roadway boundaries is associated with a systematic shift in extreme behavior and a concentration of the distribution. Relative distance shows a negative association with the location parameter, whereas its contribution to the scale component is not statistically significant at the pooled level, suggesting that reduced clearance robustly shifts the distribution toward more critical extremes across intersections, whereas its influence on dispersion is more site-dependent and uncertain. Lateral control indicators show a consistent coupled structure: both heading and steering differences exhibit positive associations with the location parameter and negative associations with the scale parameter, indicating that greater lateral misalignment is associated with shifted extreme behavior accompanied by reduced dispersion. The left-turn indicator shows a negative association with the location parameter and a positive association with the scale parameter, suggesting that left-turn contexts shift the location component while broadening dispersion, which supports threshold-dependent risk responses when exceedance probabilities are evaluated across severity levels. In contrast, speed and throttle-related dynamics act primarily through dispersion: relative speed and relative acceleration show positive associations with the scale parameter, while their pooled associations with the location component are weak, indicating that approach-speed and acceleration behavior near infrastructure is more strongly linked to variability in extremes than to a uniform location shift.
@@
The shape parameter remains negative and significant across intersections (9(c)), supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail. This indicates a physical and behavioral ceiling on severity, constrained by roadway geometry and driver response capacity. Spatial variation in $\xi$ further underscores the influence of intersection design and infrastructure placement on the extremal properties of near-miss distributions. Taken together, the intersection-level results suggest that V–I near misses are more strongly shaped by clearance, braking, and lateral maneuvers than by speed differentials. In contrast, V–V near misses are dominated by high speeds, which yield deterministic severity, whereas V–I outcomes exhibit greater variability driven by the interplay among infrastructure geometry, driver avoidance strategies, and intersection control.